# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v1

@@ V Elastic internal scattering and the Debye-Waller fraction
Density-density coupling
We consider a single tagged (foreground) electron with coordinate $\hat{\mathbf{r}}$. The interaction between the foreground electron and the background lattice is taken to be of density-density form,
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}).$ (13)
Separating the background density into its static equilibrium expectation value and fluctuations, $\rho_{\rm bg}=\langle\rho_{\rm bg}\rangle+\delta\rho_{\rm bg}$, the static component generates a periodic potential acting on the foreground electron,
@@
where $\mathbf{G}$ are reciprocal lattice vectors. The resulting foreground Hamiltonian is therefore
@@
while the remaining part of $H_{\rm int}$ describes the dynamical coupling to background fluctuations. In this way the familiar Bloch Hamiltonian emerges directly from the same density-density interaction, without being assumed a priori.
The lattice degrees of freedom (the background) are described by a Hamiltonian $H_{\rm bg}$. The coupling between foreground and background is retained in the form
where the tagged-electron density operator is
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$, (17)
and the full background density operator is
with $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j}$ as defined previously.
@@
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom.
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations—the purpose of the present analysis is to emphasize that even within standard perturbation theory there exist elastic scattering processes that are usually excluded by more restrictive formulations.
@@
In the original Fröhlich formulation, the exact factorization Eq. (10) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as an infinitely massive reference frame. At the same time, the internal displacement exponential is truncated to linear order,
@@
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the translational zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but not with the lattice as a whole.
Elastic, quasielastic, and inelastic processes
Retaining the full operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ in the background density ensures that elastic, quasielastic, and inelastic scattering processes are treated on equal footing already at first order in the density-density interaction. At this level, the interaction transfers a definite total momentum $\hbar\mathbf{q}$ from the foreground electron to the background, independent of how that momentum is subsequently absorbed. It is therefore useful to decompose the momentum transfer as $\mathbf{q}=\mathbf{q}_{\rm el}+\mathbf{q}_{\rm in}$, where $\mathbf{q}_{\rm el}$ corresponds to center-of-mass recoil of the background and $\mathbf{q}_{\rm in}$ is taken up by its internal degrees of freedom. The latter satisfies the exact kinematic relation
$\hbar\mathbf{q}_{\rm in}=-\sum_{\mathbf{k},\lambda}\hbar\mathbf{k}\,\Delta n_{\mathbf{k}\lambda},$ (25)
where $\Delta n_{\mathbf{k}\lambda}$ denotes the change in occupation of the background normal modes. In this formulation, the distinction between elastic and inelastic scattering is determined by the background state rather than by an expansion in ionic displacements.
@@
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal average of the phonon-diagonal matrix element of the internal density operator,
@@
The corresponding elastic (coherent) scattering intensity is then given by
@@
By contrast, the total scattering probability summed over all final internal states is
@@
The elastic internal branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\sum_{\alpha}p_{n}\,\big|\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
@@
showing that a substantial fraction of momentum-relaxing scattering events can be strictly elastic internally even at finite temperature.
@@
These estimates show that internal phonon excitation constitutes only a minor branching channel for momentum relaxation in copper, even at room temperature and for momentum transfers relevant to transport. The dominant mechanism for absorbing electron momentum is thus elastic or quasi-elastic coupling to the lattice background, consistent with the discussion above.
@@ VI Experiments suggesting large elastic contributions
Traditionally, the suppression of quantum-interference effects with increasing temperature has been attributed to inelastic scattering processes that transfer energy to the environment [22, 2]. However, it is important to distinguish carefully between dephasing, as operationally defined in transport experiments, and irreversible decoherence in the strict quantum-mechanical sense.
A wide range of experiments demonstrate that genuine energy relaxation remains slow over the same temperature range in which phase coherence is strongly suppressed. In particular, mesoscopic interference experiments show a rapid decrease—and in some cases saturation—of the phase coherence time $\tau_{\phi}$ at low temperatures, even as independent probes indicate that electron-phonon energy exchange rates continue to vanish [26, 16, 35]. This separation of timescales shows that momentum and phase randomization need not be accompanied by significant energy transfer.
As observed in weak-localization experiments on mesoscopic metals, the loss of quantum interference reflects the destruction of coherent backscattering between time-reversed trajectories. This coherence is far more sensitive than the transport relaxation time and can be suppressed by elastic, time-dependent scattering with negligible energy exchange and without irreversible decoherence [22, 26].
From this perspective, dephasing reflects the extreme sensitivity of quantum interference to weak, time-dependent fluctuations of the environment. Even in the absence of thermal excitations, lattice coordinates remain quantum operators with nonvanishing variance due to zero-point motion. As a result, elastic scattering amplitudes are not unity even at $T=0$, as expressed by the Debye-Waller factor. When such quantum fluctuations are slow on electronic timescales—as occurs for long-wavelength modes in finite samples—they act as an effectively time-dependent scattering potential, suppressing interference under ensemble and time averaging. These processes are elastic on electronic energy scales and do not constitute irreversible decoherence.
@@
A particularly direct separation of momentum relaxation from energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12]. In these measurements, nonequilibrium electron distributions are injected and their subsequent momentum randomization and energy loss are probed independently. Over wide temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be much shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining essentially isoenergetic. Such a pronounced separation of timescales cannot be explained by purely inelastic phonon scattering and instead points to an effectively elastic deflection channel intrinsic to the metallic state.
Johnson-noise thermometry
Johnson-noise thermometry provides a particularly stringent probe of energy relaxation in driven metallic microstructures [36, 30]. In these experiments, a dc bias drives a substantial current while the local electronic temperature is inferred directly from the Johnson-Nyquist noise. Because the noise power is proportional to the electronic energy distribution, this technique provides a quantitative and noninvasive measure of energy flow, independent of momentum transport.
A key observation is that under bias the electrical resistivity and momentum randomization rate increase strongly, while the electronic temperature remains comparatively low and spatially nonuniform. This demonstrates that electrons can undergo many momentum-deflecting scattering events without efficiently transferring energy to the lattice. Such behavior is difficult to reconcile with a picture in which inelastic phonon scattering controls momentum relaxation, but follows naturally if the dominant scattering mechanism is elastic deflection from a dynamically fluctuating lattice background.
@@ VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental facts, including the Wiedemann-Franz law, weak localization, and quantum oscillations. These topics are not speculative; they address potential objections that might otherwise arise when elastic scattering plays a dominant microscopic role. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott-Ioffe-Regel crossover. These latter points are necessarily more exploratory and are presented as a coherent physical interpretation rather than as definitive conclusions.
@@ VII.1 Implications for the Lorenz ratio
Because elastic scattering randomizes momentum just as effectively as inelastic scattering, no contradiction arises between predominantly elastic microscopic dynamics and the well-established success of conventional transport theory.
In a degenerate metal, the Wiedemann-Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the universal Sommerfeld value
@@
While some fraction of inelastic processes must exist to relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not control the transport coefficients themselves [39, 21, 31].
If the dominant momentum-relaxing channel is elastic and only weakly dependent on energy over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
@@
and the Sommerfeld value $L_{0}$ is recovered. Rare inelastic events, with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$, serve to transfer energy from the electronic system to internal lattice degrees of freedom, but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slow.
@@ VII.2 Weak localization
Weak localization (WL) provides a particularly clear illustration of the dominance of elastic backscattering in metals at low temperature. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that return to the same point after multiple elastic scattering events [6, 22]. The very existence of a WL correction therefore implies that the dominant backscattering processes are coherent and elastic.
If backscattering were predominantly inelastic, interference between time-reversed paths would be washed out and no WL correction could develop. Instead, experiments show a well-defined suppression of the conductivity at low temperature together with a characteristic negative magnetoresistance when time-reversal symmetry is broken by an applied magnetic field. As temperature is increased, WL is gradually suppressed. Although this suppression is often attributed to “inelastic scattering,” experimentally it reflects the loss of phase coherence rather than rapid energy relaxation.
Independent probes such as hot-electron relaxation and Johnson-noise thermometry show that electron-phonon energy exchange remains slow over the same temperature range. The suppression of WL therefore signals the onset of a time-dependent scattering environment that randomizes phase while leaving electronic energy essentially unchanged. Within the present framework, WL fits naturally into the hierarchy of timescales
@@
with elastic scattering producing diffusive transport and slower temporal evolution of the lattice background destroying phase coherence without necessitating phonon creation or significant energy dissipation.
VII.3 Shubnikov-de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide an independent and equally stringent demonstration that electronic scattering in metals is predominantly elastic at low temperature [33, 3]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
@@ VII.3 Shubnikov-de Haas oscillations
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. Importantly, $\tau_{q}$ is primarily sensitive to elastic momentum-randomizing processes, while genuinely inelastic energy exchange rapidly suppresses oscillations. Any scattering event that transfers an energy comparable to $\hbar\omega_{c}$ destroys the coherence required to complete a cyclotron orbit.
The observation of quantum oscillations over broad temperature ranges and down to low magnetic fields therefore implies that electrons undergo many momentum-deflecting collisions without significant energy exchange. As in the case of weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to a time-dependent scattering environment rather than rapid inelastic backscattering. Quantum oscillations thus provide direct experimental evidence that elastic scattering dominates electron dynamics in normal metals at low temperature.
VII.4 Sound attenuation
Ultrasonic attenuation in metals provides a complementary probe of electron-lattice dynamics from the lattice side. Experimentally, sound waves propagating through a metal are strongly attenuated even in very clean samples, and this attenuation is often interpreted as evidence for inelastic electron-phonon scattering [28, 25]. However, attenuation of a coherent acoustic wave does not, by itself, imply irreversible energy dissipation.
From a microscopic perspective, attenuation measures the loss of coherence of a propagating phonon mode. This loss can arise through elastic scattering processes that redistribute acoustic momentum among phonon modes or transfer momentum to the electronic subsystem, even when the total energy of the combined electron-lattice system is conserved. In this sense, sound attenuation is directly analogous to the suppression of quantum interference in electronic transport: coherent propagation is degraded without requiring strong energy relaxation [19, 1].
Standard theories of ultrasonic attenuation distinguish regimes based on the relative timescales of the sound frequency and electronic or phononic relaxation rates. In the collisionless (Landau-Rumer) regime, attenuation arises from scattering of sound by thermal excitations and mode conversion rather than from viscous energy dissipation [20, 25]. More generally, the acoustic wave loses forward coherence through elastic momentum exchange with a dynamically fluctuating background, while energy relaxation enters only at longer timescales.
Within the present framework, elastic momentum exchange between electrons and the lattice background provides a natural mechanism for ultrasonic attenuation. Time-dependent lattice distortions scatter sound waves, redirecting acoustic momentum and reducing the amplitude of the forward-propagating mode, while leaving electronic and lattice energies largely unchanged. Energy dissipation occurs subsequently through rare inelastic processes that establish thermal equilibrium.
This interpretation is consistent with experimental observations in normal metals, where substantial ultrasonic attenuation persists even in temperature regimes where independent measurements show that electron-phonon energy transfer is strongly suppressed [28, 25]. As in weak localization and quantum oscillation phenomena, sound attenuation therefore supports a picture in which elastic scattering dominates momentum transfer, while true energy relaxation plays a secondary role.
VII.5 Planckian diffusion and linear-in-$T$ resistivity
Much of the recent literature introduces Planckian behavior as a bound on a microscopic relaxation time $\tau_{\mathrm{Pl}}\sim\hbar/k_{B}T$, often interpreted as the fastest possible rate of inelastic scattering or energy dissipation [37, 14, 13, 32]. Within this narrative, transport coefficients are obtained by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [38] and here suggests a different interpretation. The primary microscopic result is the emergence of real-space diffusion with a diffusion constant

### v2

@@ V Elastic internal scattering and the Debye-Waller fraction
Density–density coupling and emergence of the Bloch potential
We take the foreground–background interaction in density–density form,
with the tagged-electron density operator
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (15)
For the background we use the exact center-of-mass factorization
where $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}_{j}^{(0)}+\hat{\mathbf{u}}_{j}$.
Separating the background density operator into its equilibrium expectation value (taken in the background center-of-mass frame) and fluctuations,
$\rho_{\rm bg}(-\mathbf{q})=\langle\rho_{\rm bg}(-\mathbf{q})\rangle+\delta\rho_{\rm bg}(-\mathbf{q}),$ (17)
the static component produces an effective periodic potential acting on the foreground electron,
@@
where $\mathbf{G}$ are reciprocal lattice vectors.
The corresponding single-particle foreground Hamiltonian becomes
@@
which is the usual Bloch Hamiltonian. This periodic potential is not assumed a priori, but emerges as the static mean-field component of the underlying density–density coupling.
The remaining part of $H_{\rm int}$ couples the foreground electron to background density fluctuations $\delta\rho_{\rm bg}$, but with the crucial difference from the clamped-lattice formulation: the full operator $\rho_{\rm bg}(-\mathbf{q})$ retains the explicit center-of-mass translation factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, which enforces exact momentum bookkeeping in the finite, untethered crystal.
@@
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom, regardless of phonon number changes.
@@
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations [4] the present analysis shows that even within standard perturbation theory there exist elastic scattering processes that are absent when the center of mass is hidden or eliminated.
@@
In the original Fröhlich formulation, the exact factorization Eq. (11) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as a clamped reference frame. At the same time, the internal displacement exponential is truncated to linear order,
@@
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the background recoil zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but the 0 mode, the lattice recoil as a whole, is absent.
@@
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal expectation value of the internal density operator,
@@
The coherent elastic scattering intensity is
@@
The total scattering intensity summed over all final internal states is
@@
The elastic (coherent) branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
@@
the familiar Debye–Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
@@
These estimates show that, for momentum transfers relevant to transport, the overwhelming fraction of scattering events are internally elastic, in the sense that they leave the lattice in the same vibrational state. Phonon creation constitutes only a minor branching channel at room temperature, even for large-angle scattering. Momentum exchange with the lattice therefore occurs predominantly through elastic or quasi-elastic coupling to the background degrees of freedom, with internal phonon excitation representing a secondary correction.
@@ VI Experiments suggesting large elastic contributions
The suppression of quantum-interference effects with increasing temperature is commonly described in terms of inelastic scattering processes that exchange energy with environmental degrees of freedom [20, 1]. It is useful, however, to distinguish between dephasing as operationally defined in transport experiments and irreversible decoherence in the strict quantum-mechanical sense.
A broad class of experiments indicates that substantial suppression of interference can occur even when direct energy relaxation to the lattice remains weak. Mesoscopic interference measurements, for example, show a pronounced reduction— and in some cases an apparent saturation—of the phase-coherence time $\tau_{\phi}$ at low temperatures, while independent probes suggest that electron–phonon energy exchange rates continue to decrease in the same regime [23, 15, 31]. This empirical separation of timescales illustrates that phase and momentum randomization need not scale directly with the rate of substantial energy transfer to internal lattice excitations.
Weak-localization is robust in experiments, implying that the backscattering is coherent and therefore elastic. [20, 23].Elastic scattering amplitudes, including diffuse non-Bragg amplitudes, are reduced by the well known Debye–Waller factor.
@@
A particularly direct separation of momentum and energy relaxation is provided by hot-electron experiments in metals [32, 26, 10, 11]. In these measurements, nonequilibrium electron distributions are injected, and the subsequent momentum randomization and energy loss are probed independently. Over broad temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be substantially shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining close to isoenergetic on the same timescale.
This pronounced separation of timescales indicates that efficient momentum randomization does not require comparably rapid energy transfer to lattice excitations. While inelastic electron–phonon scattering ultimately governs thermalization, the observed hierarchy suggests that additional scattering channels, with weak net energy exchange on electronic scales, can play an important role in momentum relaxation within the metallic state.
@@ VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations. These considerations address potential concerns regarding the role of predominantly elastic momentum-relaxing channels. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover. The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
@@ VII.1 Implications for the Lorenz ratio
Elastic scattering randomizes momentum as effectively as inelastic scattering within standard transport formalisms. Accordingly, a predominantly elastic microscopic channel does not conflict with the established success of conventional transport theory.
In a degenerate metal, the Wiedemann–Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the Sommerfeld value
@@
Although some inelastic processes must ultimately relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not determine the leading transport coefficients themselves [35, 19, 27].
If the dominant momentum-relaxing channel is elastic and only weakly energy dependent over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
@@
and the Sommerfeld value $L_{0}$ follows in the usual way. Inelastic processes with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$ transfer energy between electrons and lattice degrees of freedom but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slower.
@@ VII.2 Weak localization
Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [6, 20]. The WL correction therefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
If scattering were strongly inelastic on the transport timescale, interference between time-reversed paths would be suppressed before a WL correction could develop. Experimentally, however, a well-defined suppression of conductivity and characteristic negative magnetoresistance are observed at low temperature. As temperature increases, the WL correction is gradually reduced. Although this reduction is commonly described in terms of increasing “inelastic scattering,” what is operationally measured is the loss of phase coherence rather than a rapid increase in electronic energy relaxation.
Independent probes, including hot-electron relaxation and Johnson-noise thermometry, indicate that electron–phonon energy exchange can remain comparatively slow in the same temperature range. The suppression of WL is therefore consistent with the onset of a time-dependent scattering environment that randomizes phase while involving only modest energy transfer on electronic scales.
Within the present framework this behavior fits naturally into the hierarchy of timescales
@@
where elastic momentum randomization produces diffusive transport, while slower temporal evolution of the lattice background reduces phase coherence without requiring substantial phonon excitation or rapid energy dissipation.
VII.3 Shubnikov–de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov–de Haas and de Haas–van Alphen effects provide an independent probe of electronic scattering in metals at low temperature [29, 2]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
@@ VII.3 Shubnikov–de Haas oscillations
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. This lifetime is sensitive to momentum-randomizing processes that broaden Landau levels, while sufficiently strong inelastic energy exchange suppresses the coherence required to complete cyclotron motion. Scattering events that transfer energies comparable to $\hbar\omega_{c}$ disrupt the phase coherence of the orbit.
The observation of well-defined quantum oscillations over broad temperature ranges and down to relatively low magnetic fields therefore indicates that electrons can undergo many momentum-deflecting collisions while maintaining phase coherence over multiple cyclotron periods. As in weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to temporal fluctuations of the scattering environment. These considerations are consistent with a regime in which momentum randomization can occur without rapid or large energy exchange on electronic scales.
VII.4 Planckian diffusion and linear-in-$T$ resistivity
Much recent literature discusses Planckian behavior in terms of a microscopic relaxation time
$\tau_{\mathrm{Pl}}\sim\frac{\hbar}{k_{B}T},$
often interpreted as setting a characteristic scale for scattering or energy dissipation [33, 13, 12, 28]. Within this perspective, transport coefficients are commonly estimated by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [5, 34] and extended here suggests a complementary interpretation. The primary microscopic result is the emergence of real-space diffusion with diffusion constant